A fish swimming behind a luxury cruise liner gets caught in a whirlpool created by the ship’s propeller. If the fish has an angular velocity of 2.5 rad/s and the water in the whirlpool accelerates at 4.5 rad/s2, what will be the instantaneous angular velocity of the fish at the end of 5.0 seconds? 


1
Expert's answer
2021-03-17T16:56:25-0400

Find the angular velocity at the end of 5 s:


"\omega_f=\omega_i+\alpha t=2.5+4.5·5=25\text{ rad/s}."